OpenStudy (anonymous):

what is the x intercept of e^ (-x^2)

OpenStudy (anonymous):

There is no x intercept for that function, because no power of e equals 0.

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

wat is the derivative of the funtion?

OpenStudy (anonymous):

Use the chain rule. \[\large \frac{d}{dx}[e^{-x^2}] = e^{-x^2}\cdot \frac{d}{dx}[-x^2]\]

OpenStudy (anonymous):

-2xe^-x^2 is derivative